You have 200 feet of fencing to enclose a rectangular plot that borders on a river. If you do not fence the side along the river, find the length and width of the plot that will maximize the area. What is the largest area that can be enclosed?
Quadratic functions

Answer by Cromlix(4381)

Width = x feet
Length = 200 - 2x
Area = x(200 - 2x)
Area(x)= 200x - 2x^2
Area'(x) = 200 - 4x
A(x) = 0
200 - 4x = 0
-4x = -200
x = 50
Using Nature Table:
................ - 50 +
200 - 4x... + 0 -
So, x = 50 feet is a maximum.
Width = 50 feet
Length = 100 feet
Area = 50*100 = 5000 feet.
Hope this helps.
:-)